OpenStudy (anonymous):

What is square root 34 approximated to three decimal places?

OpenStudy (anonymous):

Are you allowed a calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but not sure how to do it

OpenStudy (anonymous):

okay what is the nearest thing you can square?

OpenStudy (anonymous):

I would say 6^2 = 36 and now we know that it must be lower then 6 and we know 5^2=25 so it must be in between 5 and 6. Now you just narrow down your answer, or you could use the square root function on your calculator.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Type in 34 on your calculator, and click \(\boxed{\sqrt{\ \ }}\) or \(\boxed{\text{sqrt}}\).